Shift summary for Owl 11/9/99 (fwd)

ugur@cuhep.Colorado.EDU
Tue, 9 Nov 1999 06:37:27 -0700 (MST)

Owl Shift, Nov. 9 1999

On Shift: S. Ledovskoy, U. Erkarslan

Hours of beam available at all intensities: 7.0 hours
Hours of beam available at nominal intensity: 7.0 hours
Hours of E799 data (special runs not included): 4.8 hours
Hours of special runs (laser scans, muon runs...): 0.3 hours

Accounting of lost time:Tape changing, Etot power supply swap,
pipeline errors.

Tape change - 14 min
Etot Power supply failure - 1 h 39 min
Pipeline errors - 20 min

========================================================================
Runs ended this shift:

14930 BEAM_799 9 million events
14931 BEAM_799 2.78 million events
14932 MUON_RUN
=======================================================================
Tapes taken to FCC this shift

none

========================================================================
Accesses during shift:
Etot power supply is replaced

========================================================================
Overview of shift: E799 datataking, Needed access to swap Etot PS,
Special Muon run (magnet settings in 799 mode), pipeline problems

========================================================================

Description of Shift:

14930 01:45 online split request

01:05 spill 199-202 livtime=0.1 error, got busy from laser
crate.
ktevana error. last permit in stream5.
pipeline error

01:26 Tapes are full. Tapes changing

14931 01:39 start new run.
There is a massage on blackboard: "Update FERA ped's"
However, FERA peds were updated ~6 hour ago (pp 193)
Sasha assumes peds were updated due to this request
but massege wasn't erased.
We did not update FERA peds before this run and we
decided to keep the massege on the black board until
future investigation

02:52 spill 54. LIVE-TIME and DAQDEAD errors. Hcc
illumination looks bad.
SDAQ red button:ENGTOT_WW
Paged Sydney and Jason. They are on their way to
address this power supply problem.

14932 03:29 Stop BEAM_799 run and started MUON_4e12 run. this is
special muon run with all magnet set to BEAM_799 mode.
We didn't change prescales on B01 and rates are very
low (1,000/min)

03:50 Sydney an Jason arrived. They are on access to swap
up ETOT power supply

04:20 Access is finished. Stop muon run.

14932 04:31 Start BEAM_799 run

05:05 Onsiplit request.

05:24 Beam quality gets worse. Call MCR

06:15 Although, beam quality/intensity was changing on
spill-by-spill basis during last 45 min, the quality
of the beam did not improve and intenstiy was <8.3E12
call MCR

06:30 Spill 90 CsI 1137 dissapears from laser illumination.
Pipeline crate7 gives error during pipeline init.
Reboot crate7. Init pipeline+kfrend.

06:50 spill 105 back to normal runnig
========================================================================
Follow up Notes on Previous Shift

========================================================================
To do on next shift

Keep an eye on CsI 1096. PMT gain is drifting slowly. See instructions
from Julie in KTeV logbook p195